40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 6 20:54:54 2012, Masha, Update, PEM, Current PEM status 
    Reply  Fri Jul 6 22:30:14 2012, Masha, Update, PEM, Current PEM status MashaDV.png
       Reply  Sat Jul 7 15:48:00 2012, Masha, Update, PEM, Current PEM status 
Message ID: 6934     Entry time: Sat Jul 7 15:48:00 2012     In reply to: 6933
Author: Masha 
Type: Update 
Category: PEM 
Subject: Current PEM status 

Quote:

Quote:

Hi everybody,

Last night I (with the help of Jenne and Jenne's advice - not to implicate her in this or anything) changed the filters for GUR1, GUR2, and STS in C1:PEM-RMS, adding a butterworth bandpass filter at each corresponding frequency band as well as a gain to convert from counts to micros/sec, and then adding a low pass filter in case of aliasing upon squaring.

Currently the seismic signals are going crazy, and producing "Nan" output on the strip graph (which leads to the instantaneously sharp spikes - which leads to the entire signal being filled on the visualizer on the wall). I checked the DataViewer output and the tdsdata output using both grep and wc, and it seems that both every single signal point is present and is a real number (also not a small real number, thereby debunking floating-point error). I'm currently not sure why seismic-strip reads out 'Nan' - perhaps because it's taking the log of 0, taking a negative log, taking the root of a negative number, or dividing by zero.

Does anyone know if the seismic-strip Nan issue is a program bug? If it's not (and therefore a filter bug), please let me know as well.

I'll be in lab for the rest of the night changing the butterworth filters to odd-order elliptic filters (at Rana's suggestion), as well as changing the cut-off frequency for the low-pass filters.

I'll E-log about it when I'm done.

Just to be sure that my numbers are correct - The STS, GUR1, and GUR2 channels all have gain 10, right? (I parsed through the e-log, and these seem to be the most recent numbers

Thanks for your help,

Masha

 

UPDATE: I changed all of the GUR1Z channels to order-5 elliptic filters. I approximated the attenuation for each one by setting the integral from _CutoffFrequency to 10^3 Hz of 10^(-Percent(f)/20) df to 0.01, where Percent(f) is a linear approximation of the relationship between the log of the frequency and the dB level (with the attenuation defining one of the points). Right now the Nan problem continues to persist, even after I loaded the coefficients. In Dataviewer, the channels look relatively normal for the past 10 minutes, as does the data when viewed with tdsdata.

 

 FIGURED IT OUT - THERE WAS A PROBLEM WITH THE LOW PASS FILTERS (TOO HIGH ORDER). FIXING IT NOW, SHOULD BE GOOD IN AN HOUR. 

ELOG V3.1.3-